OpenStudy (anonymous):

What is the slope-intercept form of the function that contains the points (5, 6) and (9, 14)?

OpenStudy (mayankdevnani):

can you find out SLOPE using these two points ?? @adjoa15

OpenStudy (anonymous):

NO CAN YOU HELP ME

OpenStudy (mayankdevnani):

if two points are given,lets say \[\large \bf (x_1,y_1)~and~(x_2,y_2)\]

OpenStudy (mayankdevnani):

then to find slope between them,we use this formula :- \[\large \bf slope(m)=\frac{y_2-y_1}{x_2-x_1}\]

OpenStudy (mayankdevnani):

first,find out slope between (5,6) and (9,14) ?? @adjoa15

OpenStudy (mayankdevnani):

using the above formula

OpenStudy (mayankdevnani):

@adjoa15 find slope using the above formula

OpenStudy (mayankdevnani):

any difficulty you are facing right now,REPLY !

OpenStudy (campbell_st):

start by drawing a diagram... it always helps 

OpenStudy (campbell_st):

when you get the slope use the point slope form \[y - y_{1} = m(x - x_{1})\] you should be able to find the slope from the diagram then use either point (5, 6) or (9, 14) and substitute then for x1 and y1 hope it helps